OpenStudy (anonymous):

What is the x-intercept of the line with the equation? -2x + 1/2y = 18 ( blank , 0)

OpenStudy (anonymous):

for x intercept put y=0/

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hey can you plzz help me

OpenStudy (anonymous):

put y=o -2x+1/2(y)=18

OpenStudy (anonymous):

-2x+1/2(0)=18

OpenStudy (anonymous):

-2x=18

OpenStudy (anonymous):

ok thx